Determine the [OH-] of a solution that is 0.230 M in HCO3-. Also determine the pH of this solution.
Answer #1
Since HCO3- is a strong acid/base, we can instantly determine the pH of it.

-Log(Molarity given) = pH
-Log(.230M)=pH
pH= .63827...
Which rounds to .638.

To find the pOH, subtract the pH from 14.
14-.638 = 13.36.

To find the concentration of [OH-], simply do 10^-pOH.
10^(-13.36) = 4.37*10^-14 (after rounding).
answered by: korain
Answer #2

H2CO3 is in equilibrium with HCO3 + H+

Ka1= 4.45×10−7

HCO3 is in equilibrium with CO32− + H+

Ka2=4.69×10−11

[H+]=√ka1*ka2

=4.57x10-9

∴pH=-log[H+]

=8.34

pOH=14-8.34=5.66

-log[OH-]=5.66

[OH-]=2.18x10-6 M

Read up more here: http://www.chembuddy.com/?left=pH-calculation&right=pH-amphiprotic-salt

answered by: abubakar